OpenStudy (anonymous):

solve x x(x-3)+1=-55/x

OpenStudy (anonymous):

I got to x^2 - 3x=-55/x -1

OpenStudy (anonymous):

thanks for viewing

OpenStudy (callisto):

Is the question \(x(x-3)+1=\frac{-55}{x}\)?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

So, first multiply both sides by x \[x[x(x-3)+1]=x(\frac{-55}{x})\]\[x^2(x-3)+1(x)=-55\]Simplify it. \[x^3-3x^2+x=-55\]\[x^3-3x^2+x+55=0\]Well... you need to solve this :|

OpenStudy (anonymous):

the original equation is x^3 - 3x^2 +x +55=0

OpenStudy (callisto):

Hmm... Sorry :S

OpenStudy (anonymous):

without help of calculator doing every thing for you.

OpenStudy (anonymous):

it's ok.

OpenStudy (anonymous):

hhh see this http://en.wikipedia.org/wiki/Cubic_function

OpenStudy (anonymous):

help me

OpenStudy (anonymous):

why

OpenStudy (anonymous):

@Callisto The equation is \[\large x^3−3x^2+x+55=0 \] ?

OpenStudy (callisto):

Yup.

OpenStudy (anonymous):

I would only know how to solve this with the newton iteration method.

OpenStudy (swissgirl):

The guy is offline neways

OpenStudy (anonymous):

I somehow doubt that they require a solution to the question given, except the topic is algorithms.

OpenStudy (callisto):

I don't know how to solve these problems when factor theorem doesn't work, at least I haven't learnt to solve this yet...

OpenStudy (anonymous):

There isn't a simple way, besides algorithms. Newton method (easy to derive by the way) \[\Large x_{n+1}=x_n- \frac{f(x_n)}{f'(x_n)} \]

OpenStudy (anonymous):

\[ \Large x_{n+1}= \frac{x(3x^2-6x+1)-(x^3-3x^2+x+55)}{3x^2-6x+1}\] So \[\Large = \frac{2x^3-3x^2-55}{3x^2-6x+1} \]

OpenStudy (anonymous):

\[f(-1)=6 \\f(6)=-4.10344 \\ f(-4.10344)=-3.20093 \\ f(-3.20093)=-2.97957 \]

OpenStudy (anonymous):

\[f(-2.97057)=-2.95607 \] Guess Wolfram doesn't approach it closer.

OpenStudy (callisto):

Wolf gives something.... complicated.

OpenStudy (anonymous):

I see identical answers on wolf and here.

OpenStudy (anonymous):

ah I think I know what you mean now, when wolf computes the exact answer it will look different of course.

OpenStudy (callisto):

You're right :| I think it takes long time for me to learn this new concept. Thanks for helping :)

OpenStudy (anonymous):

You're welcome! If you want to learn more about it, you can easily find a lot of explanations all through the internet, it's one of the most common algorithms for approaching solutions. It's actually a pretty straightforward concept, unlike the solution for cubic equations which also have a formula (-: